A rectangular coil 21cm  by 37cm  has 100 turns. This coil produces a maximum emf of 69V  when it rotates with an angular speed of 200rad/s  in a magnetic field of strength B.


Find the value of B.         
Answer #1

induced emf e = NAwB

69 = 100 * 0.21*0.37 * 200*B

B = 69/(1554)

magnetic field B = 0.0444 T
Answer #2

V = N*d(BA)/dt = NB*dA/dt

Max value of dA/dt = wA

Vmax = NBwA ==> B = Vmax/(NwA)

B = 69 / 100 * 200 * 0.0777

= 69/ 1554

= 0.044 Wb